# Disentangling Recognition and Decision Regrets in Image-Based Reinforcement Learning

### v1

@@ Abstract
In image-based reinforcement learning (RL), policies usually operate in two steps: first extracting lower-dimensional features from raw images (the “recognition” step), and then taking actions based on the extracted features (the “decision” step). Extracting features that are spuriously correlated with performance or irrelevant for decision-making can lead to poor generalization performance, known as observational overfitting in image-based RL. In such cases, it can be hard to quantify how much of the error can be attributed to poor feature extraction vs. poor decision-making. In order to disentangle the two sources of error, we introduce the notions of recognition regret and decision regret. Using these notions, we characterize and disambiguate the two distinct causes behind observational overfitting: over-specific representations, which include features that are not needed for optimal decision-making (leading to high decision regret), vs. under-specific representations, which only include a limited set of features that were spuriously correlated with performance during training (leading to high recognition regret). Finally, we provide illustrative examples of observational overfitting due to both over-specific and under-specific representations in maze environments as well as the Atari game Pong.
@@ 1 Introduction
Figure 1: Block diagram of an RL agent consisting of a recognition policy (i.e. encoder) followed by a decision policy (i.e. output head).
@@ 2 Related Work
In RL, algorithms for training policies in the form of neural networks generally fall into one of two categories: action-value fitting methods such as Mnih et al. (2015) and policy gradient methods such as Mnih et al. (2016); Schulman et al. (2017). When these neural networks take images as input, one of the most common architectures used in practice is an encoder consisting of convolutional layers followed by either an action-value head if using an action-value fitting method (e.g. Mnih et al., 2015; Raffin et al., 2021), or followed by a policy head for generating actions and a value head for estimating values if using a policy gradient method (e.g. Cobbe et al., 2019; Raffin et al., 2021). With this type of architecture (an encoder followed by output heads), the intermediary output of the encoder can be viewed as a lower-dimensional representation of the input images. While a lot of previous work has focused on learning “better” representations for various purposes (e.g. Sonar et al., 2021; Raileanu and Fergus, 2021; Dabney et al., 2021; Agarwal et al., 2021; Stooke et al., 2021; Zhang et al., 2022; Eysenbach et al., 2022), our focus in this paper is to quantify the quality of an encoder’s representations as far as its contribution to the end performance is concerned. We capture the performance loss due to a poor encoder (i.e. a poor recognition policy) through recognition regret. Meanwhile, any additional losses that might be incurred by the subsequent action-value/policy/value heads (i.e. the decision policies) are captured through decision regret.
Others have also asked what constitutes a “good” representation in RL: According to Bellemare et al. (2019); Le Lan et al. (2022), representations should be sufficient in approximating the value of all stationary policies. Meanwhile, Ghosh and Bellemare (2020); Wang et al. (2024) survey different representation learning schemes and evaluate a variety of properties such as stability/robustness, capacity, and redundancy. Different from all this work, we consider the representations extracted by a recognition policy not in isolation but in relation to a specific decision policies, which may have been trained jointly with the recognition policy or subsequently (i.e. recognition regret w.r.t. decision regret). According to our work, representations should not only limit performance (cf. recognition regret) but also facilitate the training of high-performing decision policies (cf. decision regret). Under-specific representations fail at the former while over-specific representations fail at the latter.
@@ 3 Defining Recognition and Decision Regrets
$R_{\varepsilon}(\rho_{0},\pi_{0})$ $\doteq V_{\varepsilon}^{*}-V_{\varepsilon}(\rho_{0},\pi_{0})$ (2)
$\doteq\max\nolimits_{\rho,\pi}V_{\varepsilon}(\rho,\pi)-V_{\varepsilon}(\rho_{0},\pi_{0})$ (3)
@@ 3.1 A Worked Example
Figure 2: States (left), observations (center), and representations (right) for our 3-state worked example (at time step $t=14$).
@@ 3.2 Empirical Examples
We consider two image-based environments: Minigrid, which is a maze environment where the agent needs to navigate around walls to reach some target (Chevalier-Boisvert et al., 2023), and Pong, which is an Atari game where the agent aims to score points against a computer opponent (Towers et al., 2023). In both environments, we consider policies with the same type of architecture, where the recognition policy, $z=\rho_{\text{CNN}}(x)$, is given by a convolutional neural network (CNN) and the decision policy, $a\sim\pi_{\text{MLP}}(z)$, is given by a multi-layer perceptron (MLP), and we train these policies using the PPO algorithm (Schulman et al., 2017) from the skrl package (Serrano-Muñoz et al., 2023). When presenting results, we shift and scale rewards so that the regret is at most one. Details regarding the setup can be found in the appendix.
(a) Minigrid
(b) Pong
Figure 3: Recognition and Decision Regrets under Randomized Actions and Masked Observations. Both scenarios lead to regret. For random actions, this is mostly decision regret, indicating that: Even when forced to occasionally take random actions, agents are able to learn useful representations of their environment (to be able to optimize their non-random actions). For masked observations, this is almost entirely recognition regret, indicating that: Partially masked observations fundamentally limit the ability of the agent to optimize their actions regardless of their decision policy.
@@
where $p\in[0,1]$ is the probability of taking a forced random action and $\text{Uniform}(A)$ is the uniform distribution over the action space $A$. Then, we train the weights of the two networks, $\rho_{\text{CNN}}$ and $\pi_{\text{MLP}}$, as usual via PPO.
@@
In practice, it is not always possible to directly measure the decision, recognition, or overall regret of policies as these metrics require knowing optimal policies under various constraints. Often, our goal is to find such optimal policies in the first place. For instance, estimating $\max_{\pi}V(\rho_{0},\pi)$ would require us to optimize a decision policy given a recognition policy. Suppose we happen to have a method to do so with the result being $\hat{\pi}^{*}\approx\operatorname{argmax}_{\pi}V(\rho_{0},\pi)$ so that we can estimate the decision regret as $\smash{R^{\text{dec}}}\approx V(\rho_{0},\hat{\pi}^{*})-V(\rho_{0},\pi_{0})$, which is indeed what we did in our empirical examples in Section 3.2. But in a practical application, the decision policy we are interested in analyzing, which we have been denoting as $\pi_{0}$, would most likely be the same policy as $\hat{\pi}^{*}$ in the first place, rendering our estimate meaningless as it would be equal to zero identically.
However, when measuring generalization performance, that is the performance of policies optimized in some training environment but evaluated in some other test environment, then it can be possible to estimate recognition and decision regrets in a meaningful way (essentially by computing regret via policies optimized in the test environment to analyze policies that are optimized in the training environment). Next, we follow this line of thought as a practical application of our definitions in Section 3.
@@
Instead of a single environment $\varepsilon$, we now consider two different environments: a training environment $\varepsilon^{\text{train}}$ and a test environment $\varepsilon^{\text{test}}$, which we denote together as $\mathcal{E}=(\varepsilon^{\text{train}},\varepsilon^{\text{test}})$. For instance, if we are interested in observational overfitting, these two environments might differ in terms of their observation dynamics $\omega$ (but not $\tau$). Moreover, suppose that we are given a learning algorithm $\mathcal{L}$ for training policies such that $\mathcal{L}(\varepsilon)\doteq(\mathcal{L}_{\rho}(\varepsilon),\mathcal{L}_{\pi}(\varepsilon))\approx\operatorname{argmax}_{\rho,\pi}V_{\varepsilon}(\rho,\pi)$ and $\mathcal{L}_{\pi}(\varepsilon|\rho_{0})\approx\operatorname{argmax}_{\pi}V_{\varepsilon}(\rho_{0},\pi)$. Then, the generalization performance of policies trained in $\varepsilon^{\text{train}}$ via $\mathcal{L}$ but tested in $\varepsilon^{\text{test}}$ is given by $V_{\varepsilon^{\text{test}}}(\mathcal{L}(\varepsilon^{\text{train}}))$. Similar to regret in Section 3, we can consider this performance relative to the best possible performance in the test environment, which we call the generalization regret:
@@
Estimating this metric still requires knowing the optimal test policies $\operatorname{argmax}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)$. However, unlike the case with regular regret, using the policies $\mathcal{L}(\varepsilon^{\text{test}})$ trained directly in the test environment as a proxy of the optimal test policies now leads to a meaningful estimate of $G\!R$ (since these policies are now distinct from the policies $\mathcal{L}(\varepsilon^{\text{train}})$ trained originally). This results in the following estimate:
@@
which measures how much the test performance could have been improved if it were to be possible to train policies directly in the test environment. Notice the subtle difference between $\smash{\widehat{G\!R}}$ and $G\!R$: The latter measures how much of the test performance is missed out by training policies only in the training environment. When the policies trained in the training environment are already generalizable to the test environment, that is when $G\!R=0$, for sensible algorithms $\mathcal{L}$, $\smash{\widehat{G\!R}}$ should also be zero, meaning no performance gain should be possible by re-training policies in the test environment.
@@
which measures how much performance is retained when the policies trained in the training environment are rolled out in the test environment. Both definitions are ultimately concerned with the same generalization performance given by $V_{\varepsilon^{\text{test}}}(\mathcal{L}(\varepsilon^{\text{train}}))$. The difference is that, while $G\!E$ measures this performance relative to the training performance achieved in the training environment, $\smash{\widehat{G\!R}}$ measures it relative to the training performance that could have been achieved in the test environment. However, our definition $\smash{\widehat{G\!R}}$, being a type of regret, lends itself to the same decomposition in Section 3 into recognition and decision components, which is not immediately possible with $G\!E$.
@@
Similar to Section 3, consider the intermediary case when only the decision policy has access to the test environment and the test value that would have been achieved then—that is
@@
We decompose $\smash{\widehat{G\!R}}$ with respect to this intermediary:
@@
Intuitively, $\smash{\widehat{G\!R}}{}^{\text{dec}}$ is the hypothetical performance gain from re-training $\mathcal{L}_{\pi}(\varepsilon^{\text{train}})$ in the test environment, and $\smash{\widehat{G\!R}}{}^{\text{rec}}$ is the additional performance gain from re-training $\mathcal{L}_{\rho}(\varepsilon^{\text{train}})$ as well (note that $\smash{\widehat{G\!R}}=\smash{\widehat{G\!R}}{}^{\text{rec}}+\smash{\widehat{G\!R}}{}^{\text{dec}}$). A worked example, similar to the one in Section 3.1, can be found in the appendix.
@@
  • We say that representations generated by the recognition policy $\rho_{0}$ are under-specific if $\smash{\widehat{G\!R}{}^{\text{rec}}}$ is high but $\smash{\widehat{G\!R}{}^{\text{dec}}}$ is relatively low. Intuitively, under-specific representations include “too few” features which do not contain enough information to perform the task at hand (for instance, a blind recognition policy that outputs the same representation for all observations, $\rho^{\prime}(x_{t})=\emptyset$). This limits the performance of all decision policies, leading to high $\smash{\widehat{G\!R}{}^{\text{rec}}}$.
  • We say that representations generated by the recognition policy $\rho_{0}$ are over-specific if $\smash{\widehat{G\!R}{}^{\text{dec}}}$ is high but $\smash{\widehat{G\!R}{}^{\text{rec}}}$ is relatively low. Intuitively, over-specific representations include “too many” features which still contain enough information to perform the task at hand but also contain irrelevant information. This does not limit the performance that can be achieved by subsequent decision policies, leading to low $\smash{\widehat{G\!R}{}^{\text{rec}}}$. However, it might cause a decision policy to take different actions in situations that are semantically the same (i.e. the same state $s_{t}$) just because the irrelevant information differs (i.e. different representations $z_{t}$), leading to high $\smash{\widehat{G\!R}{}^{\text{dec}}}$. [^1] Here, it is $\smash{\widehat{G\!R}{}^{\text{dec}}}$ that is higher rather than $\smash{\widehat{G\!R}{}^{\text{rec}}}$ because it is the decision policy where the overfitting happens and where the drop in performance is introduced. Still, this is the consequence of over-specific representations, generated by an underfitted recognition policy, as evidenced by the fact that a lower-regret decision policy can be found if the training algorithm is run in the test environment.
@@ 5 Illustrative Examples
In this section, we give concrete examples of observational overfitting due to both under-specific representations and over-specific representations in two image-based environments: Minigrid in Section 5.1, which is a collection of maze environments, and the Atari game Pong in Section 5.2. These environments contrast each other well in terms of their scale. [^2] Besides the scale contrast, these environments are commonly considered when investigating generalization in RL (e.g. Sonar et al., 2021; Taiga et al., 2023). Moreover, they offer organically-generated images whereas work on observational overfitting often only considers synthetically-generated images—including Song et al. (2020), which first introduced the concept. Minigrid outputs $7$-by-$7$ images with three channels corresponding to object types, object states, and color identifiers. Meanwhile, Pong, after some post-processing, outputs $164$-by-$82$ images that are grayscale and we stack the four most recent images to account for memoryless agents. As in Section 3.2, we train agents using the PPO algorithm. In all our experiments, the agents consist of a CNN that acts as the recognition policy followed by an MLP that acts as the decision policy (and followed by another parallel MLP that acts as the value head for the PPO algorithm). Details regarding the training setup can be found in the appendix.
5.1 Examples in Maze Environments
@@
Figure 4: Possible Maze Configurations in the Training and Test Environments. The agent is the red triangle; the goal is the green square. For training, the key-door color is determined based on the door location, but for testing, the key-door color is always gray (making it both an irrelevant and a spurious feature).

### v2

@@ Abstract
In image-based reinforcement learning (RL), policies usually operate in two steps: first extracting lower-dimensional features from raw images (the “recognition” step), and then taking actions based on the extracted features (the “decision” step). Extracting features that are spuriously correlated with performance or irrelevant for decision-making can lead to poor generalization performance, known as observational overfitting in image-based RL. In such cases, it can be hard to quantify how much of the error can be attributed to poor feature extraction vs. poor decision-making. To disentangle the two sources of error, we introduce the notions of recognition regret and decision regret. Using these notions, we characterize and disambiguate the two distinct causes behind observational overfitting: over-specific representations, which include features that are not needed for optimal decision-making (leading to high decision regret), vs. under-specific representations, which only include a limited set of features that were spuriously correlated with performance during training (leading to high recognition regret). Finally, we provide illustrative examples of observational overfitting due to both over-specific and under-specific representations in maze environments and the Atari game Pong.
@@ 1 Introduction
Figure 1: Block diagram of an agent consisting of a recognition policy (encoder) followed by a decision policy (output head).
@@ 2 Related Work
In RL, algorithms for training policies in the form of neural networks generally fall into one of two categories: action-value fitting methods such as Mnih et al. (2015) and policy gradient methods such as Mnih et al. (2016); Schulman et al. (2017). When these neural networks take images as input, one of the most common architectures used in practice is an encoder consisting of convolutional layers followed by either an action-value head if using an action-value fitting method (e.g. Mnih et al., 2015; Raffin et al., 2021), or followed by a policy head for generating actions and a value head for estimating values if using a policy gradient method (e.g. Cobbe et al., 2019; Raffin et al., 2021). With this type of architecture (an encoder followed by output heads), the intermediary output of the encoder can be viewed as a lower-dimensional representation of the input images. While a lot of previous work has focused on learning “better” representations for various purposes (e.g. Sonar et al., 2021; Raileanu & Fergus, 2021; Dabney et al., 2021; Agarwal et al., 2021; Stooke et al., 2021; Zhang et al., 2022; Eysenbach et al., 2022), our focus in this paper is to quantify the quality of an encoder’s representations as far as its contribution to the end performance is concerned. We capture the performance loss due to a poor encoder (i.e. a poor recognition policy) through recognition regret. Meanwhile, any additional losses that might be incurred by the subsequent action-value/policy/value heads (i.e. the decision policies) are captured through decision regret.
Others have also asked what constitutes a “good” representation in RL: According to Bellemare et al. (2019); Le Lan et al. (2022), representations should be sufficient in approximating the value of all stationary policies. Meanwhile, Ghosh & Bellemare (2020); Wang et al. (2024) survey different representation learning schemes and evaluate a variety of properties such as stability/robustness, capacity, and redundancy. Different from all this work, we consider the representations extracted by a recognition policy not in isolation but in relation to a specific decision policies, which may have been trained jointly with the recognition policy or subsequently (i.e. recognition regret w.r.t. decision regret). According to our work, representations should not only limit performance (cf. recognition regret) but also facilitate the training of high-performing decision policies (cf. decision regret). Under-specific representations fail at the former while over-specific representations fail at the latter.
@@ 3 Defining Recognition and Decision Regrets
$R_{\varepsilon}(\rho_{0},\pi_{0})$ $\doteq V_{\varepsilon}^{*}-V_{\varepsilon}(\rho_{0},\pi_{0})\doteq\max\nolimits_{\rho,\pi}V_{\varepsilon}(\rho,\pi)-V_{\varepsilon}(\rho_{0},\pi_{0})$ (2)
@@ 3.1 A Worked Example
Figure 2: States (left), observations (center), and representations (right) for our 3-state example at $t\!=\!14$.
@@ 3.2 Empirical Examples
Setup
We consider two image-based environments: Minigrid, which is a maze environment where the agent needs to navigate around walls to reach some target (Chevalier-Boisvert et al., 2023), and Pong, which is an Atari game where the agent aims to score points against a computer opponent (Towers et al., 2023). In both environments, we consider policies with the same type of architecture, where the recognition policy, $z=\rho_{\text{CNN}}(x)$, is given by a convolutional neural network (CNN) and the decision policy, $a\sim\pi_{\text{MLP}}(z)$, is given by a multi-layer perceptron (MLP), and we train these policies using the PPO algorithm (Schulman et al., 2017) from the skrl package (Serrano-Muñoz et al., 2023). When presenting results, we shift and scale rewards so that the regret is at most one. Details regarding the setup can be found in the supplementary material.
@@
where $p\in[0,1]$ is the probability of taking a forced random action and $\text{Uniform}(A)$ is the uniform distribution over $A$. Then, we train the weights of the two networks, $\rho_{\text{CNN}}$ and $\pi_{\text{MLP}}$, as usual via PPO.
@@
(a) Minigrid
(b) Pong
Figure 3: Recognition and Decision Regrets under Randomized Actions and Masked Observations. Both scenarios lead to regret. For random actions, it is mostly decision regret, indicating: Even when forced to occasionally take random actions, agents still learn useful representations of their environment (to be able to optimize their non-random actions). For masked observations, it is recognition regret, indicating: Partially masked observations fundamentally limit the agent’s ability to optimize their actions regardless of their decision policy.
@@
In practice, it is not always possible to directly measure the decision, recognition, or overall regret of policies as these metrics require knowing optimal policies under various constraints. Often, our goal is to find such optimal policies in the first place. For instance, estimating $\max_{\pi}V(\rho_{0},\pi)$ would require us to optimize a decision policy given a recognition policy. Suppose we happen to have a method to do so with the result being $\hat{\pi}^{*}\approx\operatorname{argmax}_{\pi}V(\rho_{0},\pi)$. Then, we can estimate the decision regret as $\smash{R^{\text{dec}}}\approx V(\rho_{0},\hat{\pi}^{*})-V(\rho_{0},\pi_{0})$, which is indeed what we did in our empirical examples in Section 3.2. But in a practical application, the decision policy we are interested in analyzing, which we have been denoting as $\pi_{0}$, would most likely be the same policy as $\hat{\pi}^{*}$ in the first place, rendering our estimate meaningless as it would be equal to zero identically.
However, when measuring generalization performance, that is the performance of policies optimized in some training environment but evaluated in some other test environment, then it can be possible to estimate recognition and decision regrets in a meaningful way (essentially by computing regret via policies optimized in the test environment to analyze policies that are optimized in the training environment). In this section, we develop this idea as a practical application of our definitions in Section 3. First, we introduce the notion of generalization regret to quantify how much performance is missed out in the test environment due to suboptimal generalization from the training environment. Then, similar to our approach in the previous section, we decompose generalization regret into recognition and decision components. These two components allow us to characterize over-specific vs. under-specific representations, which we exemplify later in Section 5.
@@
Instead of a single environment $\varepsilon$, we now consider two different environments: a training environment $\varepsilon^{\text{train}}$ and a test environment $\varepsilon^{\text{test}}$, which we denote together as $\mathcal{E}=(\varepsilon^{\text{train}},\varepsilon^{\text{test}})$. For instance, if we are interested in observational overfitting, these two environments might differ in terms of their observation dynamics $\omega$ (but not $\tau$). Moreover, suppose that we are given a learning algorithm $\mathcal{L}$ for training policies such that $\mathcal{L}(\varepsilon)\doteq(\mathcal{L}_{\rho}(\varepsilon),\mathcal{L}_{\pi}(\varepsilon))\approx\operatorname{argmax}_{\rho,\pi}V_{\varepsilon}(\rho,\pi)$ and $\mathcal{L}_{\pi}(\varepsilon|\rho_{0})\approx\operatorname{argmax}_{\pi}V_{\varepsilon}(\rho_{0},\pi)$. Then, the generalization performance of policies trained in $\varepsilon^{\text{train}}$ via $\mathcal{L}$ but tested in $\varepsilon^{\text{test}}$ is given by $V_{\varepsilon^{\text{test}}}(\mathcal{L}(\varepsilon^{\text{train}}))$. Similar to regret in Section 3, we can consider this performance relative to the best possible test performance, which we call the generalization regret:
@@
Estimating this metric still requires knowing the optimal test policy $\operatorname{argmax}_{\rho,\pi}V_{\varepsilon^{\text{test}}}(\rho,\pi)$. However, unlike the case with regular regret, we can now obtain a meaningful estimate of $G\!R$ by using policies $\mathcal{L}(\varepsilon^{\text{test}})$ trained directly in the test environment as a proxy for the optimal test policy (since these policies are now distinctly different from the policies $\mathcal{L}(\varepsilon^{\text{train}})$ trained originally). This results in the following estimation of the generalization regret:
@@
which measures how much the test performance could have been improved if it were to be possible to train policies directly in the test environment. Notice the subtle difference between this and $G\!R$: The latter measures how much of the test performance is missed out by training policies only in the training environment. When the policies trained in the training environment are already generalizable to the test environment, that is when $G\!R=0$, for sensible algorithms $\mathcal{L}$, $\smash{\widehat{G\!R}}$ should also be zero, meaning no performance gain should be possible by re-training policies in the test environment.
@@
which measures how much performance is retained when the policies trained in the training environment are rolled out in the test environment. Both definitions are ultimately concerned with the same generalization performance given by $V_{\varepsilon^{\text{test}}}(\mathcal{L}(\varepsilon^{\text{train}}))$. The difference is that, while $G\!E$ measures this performance relative to the training performance achieved in the training environment, $\smash{\widehat{G\!R}}$ measures it relative to the training performance that could have been achieved in the test environment. However, our definition $\smash{\widehat{G\!R}}$, being a type of regret, lends itself to the same decomposition as in Section 3 into recognition and decision components, which is not immediately possible with $G\!E$.
@@
Similar to Section 3, consider the intermediary case when only the decision policy has access to the test environment and the test value that would have been achieved then, which can be written as
@@
Then, we decompose $\smash{\widehat{G\!R}}$ with respect to this intermediary value as follows:
@@
$\widehat{G\!R}{}^{\text{dec}}_{\mathcal{E}}(\mathcal{L})$ =(10)$-V_{\varepsilon^{\text{test}}}(\mathcal{L}(\varepsilon^{\text{train}}))$ ) (12)
Intuitively, $\smash{\widehat{G\!R}}{}^{\text{dec}}$ is the hypothetical performance that could be gained from re-training $\mathcal{L}_{\pi}(\varepsilon^{\text{train}})$ directly in the test environment, and $\smash{\widehat{G\!R}}{}^{\text{rec}}$ is the further performance gain of re-training $\mathcal{L}_{\rho}(\varepsilon^{\text{train}})$ in addition to $\mathcal{L}_{\pi}(\varepsilon^{\text{train}})$ (note that $\smash{\widehat{G\!R}}=\smash{\widehat{G\!R}}{}^{\text{rec}}+\smash{\widehat{G\!R}}{}^{\text{dec}}$ is the total gain). A worked example, similar to the one in Section 3.1, can be found in the supplementary material.
@@
  • We say that representations generated by the recognition policy $\rho_{0}$ are under-specific if $\smash{\widehat{G\!R}{}^{\text{rec}}}$ is high but $\smash{\widehat{G\!R}{}^{\text{dec}}}$ is relatively low. Intuitively, under-specific representations include “too few” features which do not contain enough information to perform the task at hand (for instance, a blind recognition policy that outputs the same representation for all observations, $\rho^{\prime}(x_{t})=\emptyset$, or an overfitted recognition policy that only learned to extract spurious features). This limits the performance of all decision policies, leading to high $\smash{\widehat{G\!R}{}^{\text{rec}}}$.
  • We say that representations generated by the recognition policy $\rho_{0}$ are over-specific if $\smash{\widehat{G\!R}{}^{\text{dec}}}$ is high but $\smash{\widehat{G\!R}{}^{\text{rec}}}$ is relatively low. Intuitively, over-specific representations include “too many” features which still contain enough information to perform the task at hand but also contain irrelevant information, for instance if the recognition policy has underfitted. This does not necessarily limit the performance that can be achieved by subsequent decision policies, leading to low $\smash{\widehat{G\!R}{}^{\text{rec}}}$, but it may cause a decision policy to take different actions in situations that are semantically the same (i.e. the same state $s_{t}$) just because the irrelevant information differs (i.e. different representations $z_{t}$), leading to high $\smash{\widehat{G\!R}{}^{\text{dec}}}$. Here, it is the decision policy that has overfitted, due to the over-specific representations generated by an underfitted recognition policy.
@@ 5 Illustrative Examples
In this section, we give concrete examples of observational overfitting due to both under-specific representations and over-specific representations in two image-based environments: Minigrid in Section 5.1, which is a collection of maze environments, and the Atari game Pong in Section 5.2. These environments contrast each other well in terms of their scale. [^1] Besides the scale contrast, these environments are commonly considered when exploring generalization in RL (e.g. Sonar et al., 2021; Taiga et al., 2023). Moreover, they offer organically-generated images whereas the literature on observational overfitting tends to rely mostly on synthetically-generated images, including Song et al. (2020), which introduced the concept. Minigrid outputs $7$-by-$7$ images with three channels corresponding to object types, object states, and color identifiers. Meanwhile, Pong, after some post-processing, outputs $164$-by-$82$ images that are grayscale and we stack the four most recent images to account for memoryless agents. In addition to these examples, we also highlight model selection as a potential use case of analyzing the specificity of representations in Section 5.3.
@@
Figure 4: Possible Maze Configurations. The agent is the red triangle, and the goal is the green square. For training, the key-door color is determined based on the door location, but for testing, the key-door color is always gray (making it both an irrelevant and a spurious feature).
As in Section 3.2, we train agents using the PPO algorithm. In all our experiments, the agents consist of a CNN that acts as the recognition policy followed by an MLP that acts as the decision policy (and followed by another parallel MLP that acts as the value head for the PPO algorithm). Details regarding the training setup can be found in the supplementary material.
5.1 Examples in Maze Environments